OpenStudy (anonymous):

show the steps to find the difference please. -5(x^2 + 2x – 5) – (7x^2 – 2x – 3)

OpenStudy (anonymous):

@myininaya the champ can you help

OpenStudy (anonymous):

Distribute the -5.... -5x^2 -10x +25 -7x^2 -2x-3 then put like terms together.... -12x^2 -12x +22

OpenStudy (anonymous):

then what?

OpenStudy (anonymous):

That is the difference between the two polynomials.

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

-12x^2 -12x +22

OpenStudy (anonymous):

so thats the answer? -12x^2 - 12x + 22? if so then i was way off

OpenStudy (anonymous):

That is the answer. You can't subtract the an x term from a x^2 term and you can't subtract the single numbers (22) from the x term or x^2 term unless they gave you a value for x or gave you a complete equation; as in x-2 =0 then you know x =2. make sense?

OpenStudy (anonymous):

yes, I guess its go back to the beginning of the chapter and re-read .. thank you

OpenStudy (anonymous):

You're welcome.

myininaya (myininaya):

You guys need to distribute the negative one for the second grouping of terms -(a-b-c)=-a+b+c

OpenStudy (anonymous):

That is true. Good catch @myininaya